OpenStudy (anonymous):

how much would you need to invest to get $20ˏ000 in 5 years at an annual rate of 8.5% compounded monthly

OpenStudy (campbell_st):

well you know the annula rate, divide it by 12 to get the monthly \[20000 = P(1 + \frac{\frac{8.5}{12}}{100})^{5 \times 12}\] make p the subject... and then evalaute P being the initial investment

OpenStudy (wolf1728):

8.5 % compounded monthly = (1+(.085/12))^12= (1+0.0070833333333333)^12 = 1.08839090589264 = 8.839090589264 % annually Solving the equation for principal Principal = Total / (1 + rate)^years Principal = 20,000 / (1.08839090589264)^5 Principal = 20,000 / 1.52730059786552 Principal = 13,095.00 Here's a calculator to check your work. http://www.1728.org/compint.htm